# Absence of bulk charge density wave order in the normal state of UTe2

### v2

@@
C. S. Kengle1,2    J. Vonka3    S. Francoual4    J. Chang5    P. Abbamonte2    M. Janoschek5,6    P. F. S. Rosa1    W. Simeth1,5,6 1Los Alamos National Laboratory, Los Alamos, NM 87545, USA 2Department of Physics, University of Illinois Urbana-Champaign, Urbana, Illinois, 61801, USA 3Laboratory for X-ray Nanoscience and Technologies, Paul Scherrer Institute, Villigen PSI, Switzerland 4Deutsches Elektronen-Synchrotron (DESY), D-22607 Hamburg, Germany 5Physik-Institut, Universität Zürich, Winterthurerstrasse 190, CH-8057 Zürich, Switzerland 6Laboratory for Neutron and Muon Instrumentation, Paul Scherrer Institute, Villigen PSI, Switzerland
@@
A spatially modulated superconducting state at zero magnetic field, known as pair density wave (PDW), is a tantalizing state of matter with unique properties. PDWs have been observed in diverse spin-singlet superconductors ranging from high-temperature cuprate and Fe-based superconductors to kagome and two-dimensional chalcogenide compounds. Recent scanning tunneling microscopy (STM) studies reveal that spin-triplet superconductor UTe2 hosts an unprecedented spin-triplet PDW whose wavevectors are indistinguishable from a preceding charge-density wave (CDW) order that survives to temperatures well above the superconducting critical temperature, $T_{c}$. Whether the PDW is the mother or a subordinate order remains unsettled. Here, based on a systematic search for bulk charge order above $T_{c}$ using resonant elastic X-ray scattering (REXS), we show that the structure factor of charge order previously identified by STM is absent in the bulk within the sensitivity of REXS, which implies that the reported CDW order in the normal state of UTe2 is a surface phenomenon. Our results invite two scenarios: either the density-wave orders condense simultaneously at $T_{c}$ in the bulk, in which case PDW order is likely the mother phase, or these modulations are restricted to the surface.
I Introduction
@@
Unconventional superconductors not only provide a platform for the investigation of exotic pairing mechanisms beyond electron-phonon coupling, but already underpin existing technology such as powerful superconducting magnets or have potential for future applications such as topological quantum computing with fractionalized excitations. In particular, conventional $s$-wave superconductors are topologically trivial, whereas unconventional superconductors may be topological depending on the underlying SC order parameter. For any newly-discovered unconventional superconductor, the determination of its SC order parameter therefore becomes a central question. However, due to experimental discrepancies, disorder, and intertwined orders, such as charge-density waves (CDWs), magnetism, pair-density waves (PDWs), and nematicity, this is frequently challenging.
@@
UTe2 is a recent addition to the family of unconventional superconductors [1, 2], and its SC order parameter continues to defy consensus. Though there is strong evidence that UTe2 is a spin-triplet superconductor [3, 4, 5, 6], reports of chiral, multicomponent, and topological superconductivity [7, 8, 9, 10] have been challenged [11, 12, 13, 14, 15]. Recently, scanning tunneling microscopy (STM) measurements identified that the SC order parameter, $\Delta_{q}(\mathrm{\mathbf{k}})$, is spatially modulated at three wave-vector components, $\bm{q}_{i}$ ($i=$ 1, 2, 3) which are around the Brillouin zone boundary [16]. This pair-density wave (PDW) is inherently linked to CDW order, $\rho_{q}$, which displays the same modulation components as $\Delta_{q}$ but shifted by a phase difference $\pi$ [17, 16, 18].
@@
Notably, CDW order survives to temperatures well above the superconducting critical temperature, $T_{c}\approx 1.6$ K. As discussed in Ref. [17], this intriguing result points to two possible scenarios of intertwined order: either the SC and PDW order parameters are dominant and generate subordinate charge modulations or the SC and CDW orders dominate and give rise to pair density modulations. In the first case, a superconducting gap energy of $250$ $\mu$eV coupled to a 10 $\mu$eV PDW state must induce a CDW with a much larger energy scale of order 25 meV. In the second case, the normal-state CDW coexists with superconductivity below $T_{c}$ to generate a 10 $\mu$eV PDW state at the same wavevectors. Gu et al argue that the second case is more plausible, whereas Aishwarya et al reason that a normal-state CDW is not inconsistent with a PDW mother phase because the PDW may melt into a CDW phase that survives to higher temperatures. Whether the PDW is the mother order or a subordinate order therefore remains unsettled.
@@
Though STM is a powerful surface probe that provides both phase and domain sensitive information with high accuracy [19, 20], it is not inherently sensitive to ordered states in the bulk. As a result, STM alone is unable to make conclusive statements about correlation lengths of ordered states in the bulk of materials. In fact, signatures of CDWs observed in STM are equally likely to arise from bulk or surface states, and the two forms of order may coexist in a material [21]. Therefore, with all experimental evidence for either CDW or PDW modulations in UTe2 being restricted to surface sensitive measurements, it remains unclear whether these modulations in the SC state extend into the bulk and, if so, whether the broadening of the momentum space CDW peaks above $T_{c}$ is also a bulk phenomenon. Signatures of a phase transition when CDW order vanishes are absent in transport or thermodynamic measurements [22, 23, 24], therefore emphasizing the necessity to perform a systematic search for charge-density wave signatures using bulk-sensitive methods.
@@
Here we provide the desired bulk measurements of the charge order in UTe2 via resonant elastic x-ray scattering (REXS) performed just above the superconducting transition ($T=2.2\,$ K), which is about four times lower than the highest temperature the CDW in UTe2 is reported to exist [18].
@@
Because STM identified 5 nm-sized patches of charge order, we investigated a region in reciprocal space near two of the reported wave vectors with respect to broad signals of correlation lengths smaller than 5 nm using incident X-ray energies of 4.95 keV (Te $L_{\mathrm{1}}$ edge) and 3.73 keV (U $M_{4}$ edge). Additionally, we investigated the vicinity of one of the ordering vectors on a tight grid sufficiently dense and large enough to identify resolution limited CDW-signals using incident X-ray energy of 3.73 keV. Our main result is that the structure factor of the putative CDW is absent from the bulk within our detection limits posed by state-of-the art X-ray diffraction.
II Results
II.1 Surface state observed in STM
We first revisit the nature of CDW order identified previously in STM studies. Sample volumes probed in STM are restricted to atomic length scales and extend, at best, over two atomic layers. STM is therefore not directly sensitive to modulation periods along the surface normal $\bm{n}$.
STM studies in Refs. [17, 16, 18] were performed on $(011)$ surfaces. Here, the top layer consists of Te atoms forming a 2-D orthorombic Bravais lattice with rectangular centered unit cell (cf. Supplementary Materials Text 1 and Figure S1). STM measurements below $T_{c}$ identified superstructure-modulations, as inferred from Fourier transformed STM images, and concluded the presence of three ordering vectors ($\bm{q}_{\mathrm{1}}$, $\bm{q}_{\mathrm{2}}$, and $\bm{q}_{\mathrm{3}}$) common to both CDW and PDW order. Fig. 1(a) shows these modulations schematically within the 2D Brillouin of the Te-lattice. Black circles correspond to the Fourier components of the lattice. Charge order peaks, shown in orange, form an isosceles triangle $\Delta$q1q2q3 and are complemented to a mirror-symmetric pentagon q1q2p1pq3 by the projections of q2 and q3 onto the x-axis (p and p1).
Notably, the high symmetry points Y, Y1, L2, W, and L1 at the boundary of the three-dimensional Brillouin zone of UTe2 denoted by purple spheres form a pentagon of same shape and of nearly equal size (cf. Fig. 1b). When projected onto the $(011)$ plane, the images of W, L1, and L2 (represented by purple circles in Fig. 1a), almost coincide with the wave-vectors $\bm{q}_{\mathrm{1}}$, $\bm{q}_{\mathrm{2}}$, and $\bm{q}_{\mathrm{3}}$, suggesting therefore that the wave vectors of interest lie in the same plane at the Brillouin zone boundary.
Figure 1(b) shows an overview of the three-dimensional first Brillouin zone of UTe2 wherein characteristic high-symmetry points are shown as spheres. The points W, L1, and L2, close to which the three wave-vectors were suspected, are colored purple. Here we use the labelling convention as in Ref. [16], describing the characteristic points, which in conventional orthorhombic coordinates are given by W(12,12,$(\frac{1}{2},\frac{1}{2},\frac{1}{2})$, L1 $(0.456,0.596,0)$, and L2 $(0.456,0.404,1)$. Charge order components are shown in terms of orange spheres and their location was inferred based on the assumption that they lie in the same reciprocal space layer as the points W, L1, and L2 at the Brillouin zone boundary.
Figure 1: Multi-component surface-CDW identified by means of scanning tunneling microscopy. (a) Schematic view of fast Fourier transformed STM images, as taken on (011) surfaces. The $\beta_{2}$-axis is parallel to the crystallographic (100) axis of UTe2 and $\beta_{1}$ is perpendicular to (100) and (011). Black circles indicate the Fourier components due to the lattice of Te-atoms. Large orange circles correspond to charge-ordered peaks, $\bm{q}_{1},\bm{q}_{2},\bm{q}_{3}$. p and p1 correspond to projections of $\bm{q}_{3}$ and $\bm{q}_{2}$ onto the x-axis. The purple spheres correspond to projections of high-symmetry points (denoted by the respective capital letters) within the three-dimensional Brillouin zone. (b) High-symmetry points in the first Brillouin zone of UTe2, as represented by spheres. Ordering vectors were identified in the vicinity of W, L1, and L2 (purple spheres). The orange points correspond to the location of charge-ordered peaks within the (011) plane at the Brillouin zone boundary that have the projections that are shown in (a). (c) Exemplary cut along the line WS. Shown in terms of square symbols is intensity inferred from a fast Fourier transformation of STM picture taken at $4.9$ K (cf. Fig. 2b of Ref. [18]). For the purposes in this study, we fitted the data by a superposition of exponentially decaying background and a Gaussian peak with FWHM $f_{\mathrm{STM}}$ (brown lines). The Gaussian profile is shown in orange shading.
Above the superconducting transition temperature the charge order begins to melt. Charge order persists in the form of patches which shrink upon increasing temperature. For example, at 4.8 K the patches are of the order of $5$ nm in size. Finally, the charge order in the superconducting state is field-dependent and vanishes at the same critical field, where superconductivity is suppressed, suggesting that SC and PDW order are coupled.
Fig. 1(c) shows a typical intensity $vs.$ momentum curve obtained from a Fast Fourier Transformation (FFT) of STM images taken at 4.9 K along the line connecting the points W and S, $i.e.$, between $(\frac{1}{2},\frac{1}{2},\frac{1}{2})$ and $(\frac{1}{2},\frac{1}{2},0)$ as specified in conventional orthorhombic coordinates. A pronounced peak near the W point arises from charge-density wave modulations and is characterized by a broad Gaussian profile centered at $h=0.545\,$ r.l.u. The first wave-vector, $\bm{q}_{\mathrm{1}}=(0.545,0.5,0.5)$, is thus slightly off the W point and close to the ordering vector inferred from two dimensional Fourier transformations (See Supplementary Material Text 1 for more details). Similarly, ordering vectors close to L1 and L2 can be inferred from a detailed investigation of the STM pictures presented in Ref. [16]: $\bm{q}_{\mathrm{2}}=(0.430,0.338,1.339)$ and $\bm{q}_{\mathrm{3}}=(0.430,0.662,-0.339)$, which do not match any obvious commensurate value.
@@ II.1 Surface state observed in STM
Finally, we use the STM data to infer a conservative estimation of the expected full width at half maximum (FWHM) of bulk CDW order. Consider that the peak FWHM seen in STM represents a convolution of intrinsic peak shape ($f_{0}=2/\kappa_{0}$), associated with the correlation length $\kappa_{0}$ of CDW order, and the experimental resolution of the STM instrument, ($f_{\mathrm{STM}}$). Assuming that both $f_{0}$ and $f_{\mathrm{STM}}$ describe the widths of Gaussian profiles, we find that the profile in Fig. 1(b) has a FWHM $f_{\mathrm{W,STM}}=\sqrt{f_{0}^{2}+f_{\mathrm{STM}}^{2}}$. The experimental profile therefore serves as an upper bound for the internal width of charge-order, yielding $f_{\mathrm{max}}:=f_{\mathrm{W,STM}}=0.1439\,\mathrm{\AA}^{-1}>f_{0}$. A lower bound may be obtained considering that coherent CDW correlations were observed in terms of patches of less than $\kappa_{1}=5$ nm diameter, implying a lower bound for the intrinsic peak width of $f_{\mathrm{min}}:=2/\kappa_{1}<f_{0}$, cf. Ref. [25]. Combined, these limits on the FWHM can be written $f_{\mathrm{min}}=0.04\,\mathrm{\AA}^{-1}<f_{0}<0.1439\,\mathrm{\AA}^{-1}=f_{\mathrm{W,STM}}$.
@@
II.2 Resonant X-ray scattering
To study bulk microscopic properties of charge modulations in UTe2, we utilized resonant elastic scattering of hard X-rays and searched for signatures of a multi-component CDW, as reported previously to exist at the Brillouin zone boundary. With penetration depths in UTe2 exceeding 200200200\,200nm, these experiments are sensitive to the structure factor of long-range correlations across several hundred crystallographic unit cells. Bragg scattering at charge-order peaks is typically studied with non-resonant scattering and characteristically leaves incident polarization invariant. The beam can be linearly polarized perpendicular to the scattering plane, $\sigma$, or perpendicular to the scattering plane, $\pi$.
For the putative resonant enhancement of charge-density wave order, the structure factor is unknown, but may have finite contributions in any polarization channel. Therefore, the chance to observe charged-ordered Bragg peaks may be maximized when polarization is maintained parallel to the scattering plane, i.e., using the $\sigma\sigma^{\prime}$-channel. In our experiments, structural Bragg peaks measured on UTe2 were limited by experimental resolution represented by Gaussian profiles, indicating ideal crystalline quality. Figure 2 shows representative scans through (a) the $(011)$ Bragg peak using an incident energy of $3.73$ keV and (b) the $(022)$ Bragg peak using an incident energy of $4.95$ keV. See Supplementary Information for more details on the setup of our diffraction experiments.
II.3 Survey for putative CDW order in bulk

### v3

@@ I Title
II Authors
C. S. Kengle1,2, J. Vonka3, S. Francoual4, J. Chang5, P. Abbamonte2, M. Janoschek5,6, P. F. S. Rosa1, and W. Simeth1,5,6
III Affiliations
1Los Alamos National Laboratory, Los Alamos, NM 87545, USA
2Department of Physics, University of Illinois Urbana-Champaign, Urbana, Illinois, 61801, USA
3Laboratory for X-ray Nanoscience and Technologies, Paul Scherrer Institute, Villigen PSI, Switzerland
4Deutsches Elektronen-Synchrotron (DESY), D-22607 Hamburg, Germany
5Physik-Institut, Universität Zürich, Winterthurerstrasse 190, CH-8057 Zürich, Switzerland
6Laboratory for Neutron and Muon Instrumentation, Paul Scherrer Institute, Villigen PSI, Switzerland
IV Abstract
@@
A spatially modulated superconducting state, known as pair density wave (PDW), is a tantalizing state of matter with unique properties. Recent scanning tunneling microscopy (STM) studies revealed that spin-triplet superconductor UTe2 hosts an unprecedented spin-triplet, multi-component PDW whose three wavevectors are indistinguishable from a preceding charge-density wave (CDW) order that survives to temperatures well above the superconducting critical temperature, $T_{c}$. Whether the PDW is the mother or a subordinate order remains unsettled. Here, based on a systematic search for bulk charge order above $T_{c}$ using resonant elastic X-ray scattering (REXS), we show that the structure factor of charge order previously identified by STM is absent in the bulk within the sensitivity of REXS. Our results invite two scenarios: either the density-wave orders condense simultaneously at $T_{c}$ in the bulk, in which case PDW order is likely the mother phase, or the charge modulations are restricted to the surface.
V Introduction
@@
Unconventional superconductors not only provide a platform for the investigation of exotic pairing mechanisms beyond electron-phonon coupling [1, 2, 3, 4], but already underpin existing technology such as powerful superconducting magnets or have potential for future applications such as topological quantum computing with fractionalized excitations [5]. In particular, most conventional $s$-wave superconductors are topologically trivial, whereas unconventional superconductors may be topological depending on the underlying superconducting (SC) order parameter. For any newly-discovered unconventional superconductor, the determination of its SC order parameter therefore becomes a central question. However, this task is rendered difficult due to experimental discrepancies, disorder, and intertwined orders, such as charge-density waves (CDWs), magnetism, pair-density waves (PDWs), and nematicity [6, 7].
@@
UTe2 is a recent addition to the family of unconventional superconductors [8, 9], and its SC order parameter continues to defy consensus. Though there is strong evidence that UTe2 is a spin-triplet superconductor [10, 11, 12, 13], reports of chiral, multicomponent, and topological superconductivity [14, 15, 16, 17] have been challenged [18, 19, 20, 21, 22]. Recently, scanning tunneling microscopy (STM) measurements identified that the SC order parameter, $\Delta_{\bm{q}}(\bm{R})$, is spatially modulated at three wave-vector components, $\bm{q}_{i}$ ($i=$ 1, 2, 3), which are around the Brillouin zone boundary [23] and which display different field and temperature variations [23]. This pair-density wave (PDW) is inherently linked to CDW order, $\rho_{\bm{q}}(\bm{R})$, which displays the same modulation components as $\Delta_{\bm{q}}(\bm{R})$ but shifted by a phase difference $\pi$ [24, 23, 25].
@@
Notably, CDW order survives to temperatures well above the superconducting critical temperature, $T_{c}\approx 1.6$ K. As discussed in Ref. [24], this intriguing result points to two possible scenarios of intertwined order: either the SC and PDW order parameters are dominant and generate subordinate charge modulations or the SC and CDW orders dominate and give rise to pair density modulations. In the first case, a superconducting gap energy of $250$ $\mu$eV coupled to a 10 $\mu$eV PDW state must induce a CDW with a much larger energy scale of order 25 meV. In the second case, the normal-state CDW coexists with superconductivity below $T_{c}$ to generate a 10 $\mu$eV PDW state at the same wavevectors. Gu et al [24] argue that the second case is more plausible, whereas Aishwarya et al [23] reason that a normal-state CDW is not inconsistent with a PDW mother phase because the PDW may melt into a CDW phase that survives to higher temperatures. Whether the PDW is the mother order or a subordinate order therefore remains unsettled.
@@
Though STM is a powerful surface probe that provides both phase and domain sensitive information with high accuracy [26, 27], it is not inherently sensitive to ordered states in the bulk and therefore unable to make conclusive statements on the penetration depth of a state. In fact, signatures of CDWs observed in STM are equally likely to arise from bulk or surface states, and the two forms of order may coexist in a material [28]. Therefore, with all experimental evidence for either CDW or PDW modulations in UTe2 being restricted to surface sensitive measurements, it remains unclear whether these modulations in the SC state extend into the bulk and, if so, whether the broadening of the momentum space CDW peaks above $T_{c}$ is also a bulk phenomenon. Signatures of a phase transition when CDW order vanishes are absent in transport or thermodynamic measurements [29, 30, 31], therefore emphasizing the necessity to perform a systematic search for charge-density wave signatures using microscopic bulk-sensitive methods.
@@
Here we provide the desired bulk measurements of the charge order in UTe2 via resonant elastic X-ray scattering (REXS) performed just above the superconducting transition ($T=2.2\,$ K), which is about four times lower than the highest temperature the CDW in UTe2 is reported to exist [25]. Resonant diffraction can increase intensity of charge order by more than three orders of magnitude compared to non-resonant scattering [32, 33], where intensities are frequently prohibitively weak and which is largely insensitive to weak forms of charge-order not involving lattice degrees of freedom. In particular, non-resonant scattering precludes to rule out the presence of bulk long-range order. Therefore, having used incident X-ray energies of 4.95 keV (Te $L_{\mathrm{1}}$ edge) and 3.73 keV (U $M_{4}$ edge) is central for the conclusion of our study on the absence of bulk charge order in UTe2.
@@
Because STM above the superconducting transition identified nanometer-sized patches of charge order, we investigated a region in reciprocal space near two of the reported wave vectors with respect to broad signals of correlation lengths smaller than typical patch sizes. Additionally, to consider a state that has a different correlation length in the bulk, we investigated the vicinity of one of the ordering vectors on a tight grid sufficiently dense and large enough to identify resolution limited CDW-signals using incident X-ray energy of 3.73 keV. Our main result is that in both cases the structure factor of the putative CDW is absent from the bulk within our detection limits posed by resonant X-ray diffraction.
VI Results
VI.1 Surface charge order observed in STM
We first revisit the nature of charge order identified previously in STM studies [24, 23, 25]. Data were taken on $(011)$ planes of UTe2. The top layer after cleaving consists of Te atoms that form a 2-dimensional orthorhombic Bravais lattice with a rectangular centered unit cell (cf. Supplementary Materials Text 1 and Figure S1). Fig. 1(A) illustrates schematically characteristic signatures in two-dimensional reciprocal space, as seen in Fourier transformed STM images and discussed in Refs. [24, 23, 25]. The periodic arrangement of Te-atoms results in relatively strong structural Fourier-components (black circles). Below $T_{c}$ long-range charge-order is observed with components $\mathrm{q}_{\mathrm{1}}=(0,0.57)$, $\mathrm{q}_{\mathrm{2}}=(1,0.43)$, and $\mathrm{q}_{\mathrm{3}}=(-1,0.43)$, shown in Fig. 1(A) in terms of orange circles. With STM being restricted to atomic length scales and extending, at best, over two atomic layers, these studies were, however, unable to directly identify modulation components along $(011)$.
We will now consider putative bulk charge order with wave-vectors that result in these three surface projections. Although a single wave-vector (single-$\bm{q}$ with $\bm{q}=(0,0,0.57)$) lacking a modulation along $(011)$ could on its own lead to this pattern of projections, it is unable to account for the multi-component behavior observed through field and thermal variations (cf. Ref.[23]). We will therefore consider three independent components $\bm{q}_{\mathrm{1}}$, $\bm{q}_{\mathrm{2}}$, and $\bm{q}_{\mathrm{3}}$ in the three-dimensional Brillouin zone of UTe2.
[…]
Figure 1(B) illustrates a section of the 3D Brillouin zone for UTe2. The relevant characteristic high-symmetry points and estimated CDW wave vectors are show in purple and orange, respectively. Other high-symmetry points are indicated with grey spheres. The estimated CDW wave-vector positions do not match any obvious commensurate value, but are in the plane perpendicular to $\Gamma$S ($\Gamma$ is the origin). The first wave-vector, $\bm{q}_{\mathrm{1}}=(0.57,0.5,0.5)$, is slightly off the W point, $\bm{q}_{\mathrm{2}}=(0.430,0.338,1.339)$ and $\bm{q}_{\mathrm{3}}=(0.430,0.662,-0.339)$ are close to the Brillouin zone corners L1 and L2.
[…]
Above the superconducting transition temperature the charge order begins to melt, but persists in the form of patches which shrink upon increasing temperature [34]. For example, at 4.8 K the patches are of the order of order $5$ nm in size. Finally, we use exemplary STM data above $T_{c}$ to infer a conservative estimation of the expected full width at half maximum (FWHM) of bulk CDW order. Fig. 1(C) shows a typical intensity $vs.$ momentum curve obtained from a Fast Fourier Transformation (FFT) of STM images taken at 4.9 K along the line connecting the points W and S, $i.e.$, between $(\frac{1}{2},\frac{1}{2},\frac{1}{2})$ and $(\frac{1}{2},\frac{1}{2},0)$ as specified in conventional orthorhombic coordinates. A pronounced peak near the W point arises from charge-density wave component $\bm{q}_{1}$ and is characterized by a broad Gaussian profile centered at $H=0.545\,$ r.l.u..
@@ VI.1 Surface charge order observed in STM
Consider now, that the peak FWHM seen in STM represents a convolution of intrinsic peak shape ($f_{0}=2/\kappa_{0}$), associated with the correlation length $\kappa_{0}$ of CDW order, and the experimental resolution of the STM instrument, ($f_{\mathrm{STM}}$). Assuming that both $f_{0}$ and $f_{\mathrm{STM}}$ describe the widths of Gaussian profiles, we find that the profile in Fig. 1(B) has a FWHM $f_{\mathrm{W,STM}}=\sqrt{f_{0}^{2}+f_{\mathrm{STM}}^{2}}$. The experimental profile therefore serves as an upper bound for the internal width of charge-order, yielding $f_{\mathrm{max}}:=f_{\mathrm{W,STM}}=0.1439\,\mathrm{\AA}^{-1}>f_{0}$. In turn, a lower bound may be obtained considering that coherent CDW correlations were observed in terms of patches of less than $\kappa_{1}=5$ nm diameter, implying the bound for the intrinsic peak width of $f_{\mathrm{min}}:=2/\kappa_{1}<f_{0}$, cf. Ref. [35]. Combined, these limits on the FWHM can be written $f_{\mathrm{min}}=0.04\,\mathrm{\AA}^{-1}<f_{0}<0.1439\,\mathrm{\AA}^{-1}=f_{\mathrm{max}}$.
@@
VI.2 Resonant X-ray scattering
To study bulk microscopic properties of charge modulations in UTe2, we utilized resonant elastic scattering of hard, linearly polarized X-rays and searched for signatures of a multi-component CDW around the Brillouin zone boundary. With penetration depths in UTe2 exceeding 200200200\,200nm, these experiments are sensitive to the structure factor of long-range correlations across several hundred crystallographic unit cells.
For the putative resonant enhancement of charge-density wave order, the precise structure factor is unknown. But it is well established, that in the non-resonant case when using linearly polarized X-rays diffraction intensities from charge-order may have maximum intensity in the polarization channel $\sigma\sigma^{\prime}$, i.e., with both incident and scattered polarization parallel to the scattering plane, [36]. Therefore, in order to maximize the chance to observe charged-ordered Bragg peaks we maintain incident polarization parallel to the scattering plane [36].
In our experiments, structural Bragg peaks measured on UTe2 were limited by experimental resolution represented by Gaussian profiles, indicating ideal crystalline quality of our sample. Figure 2 shows representative scans through (A) the $(011)$ Bragg peak using an incident energy of $3.73$ keV and (B) the $(022)$ Bragg peak using an incident energy of $4.95$ keV. See Supplementary Information, Text S2, for more details on the setup of our diffraction experiments.
VI.3 Survey for putative CDW order in bulk